Determine the complex number z which satisfies the equations |z + 3i| = |z + 5 − 2i| and |z − 4i| = |z + 2i| simultaneously.
1
Expert's answer
2020-05-20T18:54:13-0400

Let "z=x+iy", then "|z+3i|^2=|x+(y+3)i|^2=x^2+(y+3)^2", "|z+5-2i|^2=|(x+5)+(y-2)i|^2=(x+5)^2+(y-2)^2", "|z-4i|^2=x^2+(y-4)^2", "|z+2i|^2=x^2+(y+2)^2", so we obtain the system "\begin{cases}
x^2+(y+3)^2=(x+5)^2+(y-2)^2\\
x^2+(y-4)^2=x^2+(y+2)^2
\end{cases}"

From the second equation we obtain "(y-4)^2=(y+2)^2", that is "y^2-8y+16=y^2+4y+4", so "y=1".

Substitute "y=1" to the first equation. We obtain "x^2+16=(x+5)^2+1", that is "x^2+16=x^2+10x+26", so "x=-1".

Answer: "z=-1+i"